# Challenges and Requirements for Benchmarking Time Series Foundation Models

### v2

@@ Abstract
Time Series Foundation Models (TSFMs) represent a new paradigm for time-series forecasting, promising zero-shot predictions without the need for task-specific training or fine-tuning. However, similar to Large Language Models (LLMs), the evaluation of TSFMs is challenging: as training corpora grow increasingly large, it becomes difficult to ensure the integrity of the test sets used for benchmarking. Our investigation of existing TSFM evaluation studies identifies two kinds of information leakage: (1) train-test sample overlaps arising from the multi-purpose reuse of datasets and (2) temporal overlap of correlated train and test series. Ignoring these forms of information leakage when benchmarking TSFMs risks producing overly optimistic performance estimates that fail to generalize to real-world settings. We therefore argue for the development of novel evaluation methodologies that avoid pitfalls already observed in both LLM and classical time-series benchmarking, and we call on the research community to adopt principled approaches to safeguard the integrity of TSFM evaluation.
@@ 1 Introduction
Time Series Foundation Models (TSFMs) represent an emerging paradigm in forecasting, drawing inspiration from the architecture and training methodologies of foundation models in natural language processing (NLP). Despite being a relatively new research area, the field has evolved into a highly dynamic landscape featuring a wide variety of architectures and approaches. This rapid proliferation is exemplified by diverse models such as Chronos (Ansari et al., 2024a), Chronos-bolt (Ansari et al., 2024b), Chronos-2 (Ansari et al., 2025), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024), Time-MoE (Shi et al., 2024), TimesFM (Das et al., 2024), TimesFM-2.0 (Das et al., 2024), TimesFM-2.5 (Das et al., 2024), or TiRex (Auer et al., 2025). In contrast to traditional time series models that are trained from scratch for specific tasks and time series, TSFMs leverage transfer learning: they are pre-trained on large collections of generic time series and can perform zero-shot forecasting without domain adaptation or fine-tuning (Liang et al., 2024).
@@
The reliance on massive amounts of time series training data, typically scraped from public repositories, raises the question of whether the proper evaluation of TSFMs faces similar challenges to those observed in the evaluation of Large Language Models (LLMs). In the NLP domain, training on vast portions of the internet has given rise to an “evaluation crisis" (Liao and Xiao, 2023), in which test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) and memorization effects (Chang et al., 2024) have led to overly optimistic performance estimates. Whether current TSFM evaluations are vulnerable to analogous issues remains an open question, with important implications for fair and reliable model comparison.
While benchmark strategies such as benchmarking on held-out data (Aksu et al., 2024) and clean train-test splits (Qiu et al., 2024) have been proposed recently, we doubt that these approaches effectively prevent information leakage in the context of TSFMs. Beyond the risk of direct information leakage, the increased memorization capacity of scaled foundation models poses additional challenges (see Section 3).
This article examines the validity and reliability of current TSFM evaluation practices and discusses the risk of an evaluation crisis in the TSFM domain. We address this issue through two contributions: first, we conduct a comprehensive train-test set lineage analysis of 22 leading TSFMs and review all major evaluation benchmarks available as of early 2026. Second, we empirically demonstrate the phenomenon of temporal overlaps of correlated train and test series, a challenge specific to TSFMs, through a dedicated experiment.
Our investigation yields two primary findings that suggest that TSFM evaluation is indeed compromised:
  • First, while we find no evidence of unintended direct information leakage within the training and evaluation of individual TSFMs, our analysis reveals substantial ambiguity in data separation between models. In several cases, one model’s training set serves as another’s test set. This issue is further exacerbated by dataset remixing, renaming, and inadequate documentation, which already pose significant evaluation challenges for newer models (see Section 3.1).
  • Second, we identify the temporal overlap of correlated train and test series as an indirect form of information leakage specific to TSFMs. External shocks, such as the COVID-19 pandemic to stock prices, as well as exogenous drivers like local weather conditions to local solar production, act as factors that directly shape or distort the causal dynamics between multiple time series (see Section 3.2). As a result, a TSFM trained and evaluated on disjoint time series covering the same time period can effectively memorize patterns induced by such causal relations (see Figure 3).
Based on these findings, we argue that the rise of TSFMs necessitates new benchmarking methodologies to prevent an evaluation crisis akin to that observed in NLP.
2 Background
@@
2.1 Time Series Foundation Models
A foundation model is defined as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted or fine-tuned to a wide range of downstream tasks” (Bommasani et al., 2021). The term is primarily associated with Transformer-based Large Language Models (LLMs) like the ChatGPT or Llama families (OpenAI, 2023; Grattafiori et al., 2024). Their success is largely attributed to scaling laws (Kaplan et al., 2020), describing the practice of increasing training data, parameters, and compute to improve a model’s predictive performance. Although LLMs demonstrate impressive zero-shot capabilities across benchmarks (Liang et al., 2022), concerns persist that they partly rely on memorizing benchmark data (Chang et al., 2024; Carlini et al., 2022). Such test set contamination, where test samples appear in pre-training data like Common Crawl [^1] https://commoncrawl.org/ or The Pile [^2] https://pile.eleuther.ai/, remains common despite improved dataset curation, leading to inflated performance estimates (Li et al., 2024a; Singh et al., 2024; Grattafiori et al., 2024; Mirzadeh et al., 2024; Ravaut et al., 2024).
Recognizing the structural similarity between language modeling and time series analysis, recent research has adapted the Transformer architecture to forecast future values of time series, ranging from the classical encoder-decoder Transformer (Ansari et al., 2024a) over Decoder-only models (Das et al., 2024) to Encoder-only models often applied in multivariate time series forecasting (Goswami et al., 2024). Beyond this, TSFMs are based on a variety of other architectural backbones such as LLMs Gruver et al. (2024); Xue and Salim (2024), Vision-based Autoencoders Chen et al. (2024), and mixed architectures that leverage LLM reprogramming Jin et al. (2024). Across these approaches, a key commonality lies in how data are represented: LLM-based models encode numerical values directly as language tokens like LLMTime (Gruver et al., 2024) and PromptCast (Xue and Salim, 2024), whereas some Transformer-based variants rely on specialized tokenization schemes to discretize continuous time series, as exemplified by TimeGPT-1 (Garza et al., 2024) and the T5-based Chronos (Ansari et al., 2024a). Furthermore, Vision-based approaches represent time series as gray-scaled images, while reprogramming-based approaches mix textual and continuous time series representations Jin et al. (2024). Similar to LLMs, all presented architectures are pre-trained on large collections of generic time series and appear to follow scaling laws (Edwards et al., 2024), enabling them to produce zero-shot forecasts and handle a wide range of frequencies and forecasting horizons. Only one exception—MOMENT—requires task-adaptive fine-tuning for high-quality forecasts Goswami et al. (2024). Based on an encoder-only architecture this model is capable of time series regression, classification, anomaly detection, and forecasting Goswami et al. (2024). However, due to its broad application area, the model needs task-adaptive fine-tuning of the last neural network layer that is specialized for each task. A key distinction between the foundation models in NLP and time series analysis concerns the definition of “zero-shot” capabilities. While LLMs typically adapt to entirely different tasks (Brown et al., 2020), TSFMs typically perform zero-shot forecasting by generalizing across domains, time series, input lengths, frequencies, and horizons without the need for model retraining (Ansari et al., 2024a; Das et al., 2024; Woo et al., 2024).
2.2 Benchmarking Methodologies
To rigorously assess a model’s predictive capabilities under evolving temporal dynamics, Time-Series Cross-Validation (TSCV) is widely regarded as best practice. Instead of using a single time-based split, TSCV uses the idea of rolling windows (see left in Figure 1). This evaluation strategy does not necessarily require retraining; instead, the training set can be fixed while the inference window is rolled forward, reducing evaluation costs to inference-only computation (Hewamalage et al., 2023).
Forecasting competitions, like the M-competitions or Kaggle, try to avoid the risk of test set contamination by soliciting forecasts for unpublished private test sets (Makridakis et al., 2022; Bojer and Meldgaard, 2021). These competitions are typically held at irregular intervals.
While all of the above evaluation strategies rely on historical data for testing, the recent M6 competition introduced a live evaluation using future values of financial assets on selected dates as the target (Makridakis et al., 2024).
While current TSFM benchmarking studies also rely on historical evaluation data, their logic diverges from the above described time-series cross-validation strategy. Instead of using TSCV on one series as illustrated in Figure 1 (a), they typically evaluate models using many series but with only a single time-based split per series (Das et al., 2024; Gruver et al., 2024). Figure 1 (b) shows the focus of most TSFM evaluations. Here, the focus is not on temporal depth, but on generalization across heterogeneous series, especially in zero-shot scenarios where complete datasets or domains are held out of the pre-training (Shi et al., 2024; Ansari et al., 2024a). This approach is focused on the generalization capabilities of TSFMs and prevents cherry-picking of test sets. Two prominent examples of this strategy are TSFM-Bench, which comprises 21 test sets with time series from diverse domains (e.g., energy, finance), statistical characteristics (e.g., stationary, trend), and frequencies (e.g., quarter-hourly, hourly) Li et al. (2025) and GIFT-Eval, which comprises 23 curated test sets with time series of seven domains, ten frequencies, and multivariate inputs (Aksu et al., 2024).
While this series cross validation focuses on evaluating the generalization capabilities of TSFMs, it is not without drawbacks. Roque et al. (2025) showed that with just four cherry-picked test sets, out of the distribution of available test sets, 46% of models could appear as state-of-the-art purely due to selection bias. Notably, deep learning models were found to be more susceptible to this variance than classical baselines. Especially when evaluation general-purpose TSFMs, generalization is a core evaluation focus.
Figure 1 (c) illustrates the current best practice in TSFM evaluation, that is, a combined time and series cross validation (Yu et al., 2016; Salinas et al., 2020). While this approach has first adopters, notably Goktas et al. (2025b); Shchur et al. (2025), it is not yet commonly used in TSFM benchmarking.
The scaling of pre-training and evaluation datasets in TSFMs gives rise to another challenge that can also be observed in LLM evaluation. As LLMs are pre-trained on large datasets crawled from the internet like Common Crawl [^3] https://commoncrawl.org/ or The Pile [^4] https://pile.eleuther.ai/, they have already seen most benchmark tasks during pre-training. Ultimately, this led to an LLM “evaluation crisis” Liao and Xiao (2023). Likewise, with large and obscure pre-training datasets, defining test datasets that do not overlap with models’ pre-training datasets becomes increasingly difficult in TSFMs Aksu et al. (2024); Goktas et al. (2025a); Xu et al. (2025). As shown in the GIFT-Eval ablation studies, information leakage can strongly inflate evaluation performance of TSFMs Aksu et al. (2024). Thus, investigating information leakage in TSFMs is an important research direction. Addressing this, current benchmarks developed strategies to prevent information leakage. GIFT-Eval provides an explicit pretraining dataset containing approximately 230 billion data points and a dedicated test with 177 million data points Aksu et al. (2024) staying with a series cross validation (see Figure 1 (a)) but no time cross validation. Two other benchmarks TempusBench and Fidel-TS, use test datasets from unknown or restricted sources to prevent information leakage Goktas et al. (2025a); Xu et al. (2025). Particularly, TempusBench leverages a new test set with 48 uni- and multi-variate time series forecasting tasks that is currently not in any TSFM pre-training corpus Goktas et al. (2025a). While they apply a time and series cross-validation strategy, they further propose to leverage real-world live data for evaluations Goktas et al. (2025a). Fidel-TS apply a similar strategy by leveraging restricted live APIs but implement a series cross validation without time cross validation Xu et al. (2025).
Overall, all current approaches to mitigate information leakage and the possibly resulting inflation of performance rely on the development of new datasets that can currently not be in the pre-training corpus of TSFMs.
3 Challenges of Evaluating Time Series Foundation Models
Based on the need for an examination of information leakage in current TSFMs, we decided to first conduct a TSFM train-test dataset lineage analysis (Section 3.1). We then further analyze the outcomes of potential indirect overlaps in time series between pre-training and test sets of TSFMs.
3.1 Direct Information Leakage: Train-Test Sample Overlap
To examine potential test set contamination in existing TSFM evaluations, we searched for peer-reviewed and accepted papers including the terms "Time Series Foundation Model" at leading venues in the machine learning area (TMLR, ICML, NeurIPS, ICLR, WWW) and their workshops, published up to January 2026. This search resulted in 22 published TSFMs that met our inclusion criteria (Table 1). For each of these papers, we analyzed their dataset lineage, categorizing the used datasets into pre-training, train/test and zero-shot.
@@ 3.1 Direct Information Leakage: Train-Test Sample Overlap
Chronos Ansari et al. (2024a) 2024 TMLR
@@
Pre-training datasets are used to fit a model to general time series patterns, such as, trends and seasonality. They are employed for initial training only and, hence, should be excluded from any evaluation activities. Train/test datasets, also sometime called in-domain or in-distribution datasets in the context of TSFMs, are datasets used to fine-tune a model to domain-specific patterns. They are split at a defined point in time into training and test sets and the model is fitted on the training data and evaluated on the test data (Hyndman and Athanasopoulos, 2021; Kapoor and Narayanan, 2023). Zero-shot datasets are used for evaluation of TSFMs on unseen data and should therefore be excluded from any training and fine-tuning. In a zero-shot setting, a model is given a short historical context of an unseen time series and it is assessed how well it can predict future values of this series.
@@
Our lineage analysis, the details are described in Supplementary Information Section D and the results are visualized in Figure 2, reveals that there is no consensus between model developers which publicly available datasets should be used for what purpose. In fact, each TSFM team uses a different combination of pre-training, train/test, and zero-shot datasets.
Whereas we find no indication that any individual TSFM suffers from information leakage within its own training and evaluation, the widespread practice of re-purposing datasets introduces substantial challenges for benchmarking across TSFMs. Once a dataset has been used for pre-training or fine-tuning by any model, it must be regarded as contaminated and is no longer admissible for TSFM benchmarking. Our lineage analysis shows that this is not merely a theoretical concern, but a concrete and pervasive practical issue. No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Supplementary Information Section D). And only 6% of the analyzed datasets have never been used for pre-training or fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
In Supplementary Information Section B, we summarize literature showing that leakage between training and test sets can easily lead to over 50% better test scores. Several documented leakage cases suggest that the often unintended overlaps between train and test data are not always identified during the peer-review process Saravanan et al. (2024); Montet et al. (2025); Li et al. (2024b). This is not surprising as exactly determining on which datasets a model has been trained requires careful reading of the original paper, its appendix, and sometimes even analyzing published source code. Doing this for a dozen or more models when performing a benchmarking study requires substantial effort from the authors and reviewers as well. As more TSFMs are published, authors increasingly face the challenge of navigating complicated dataset relationships. For example, the authors of the Tirex model discovered that a dataset called "Solar" appears in parts in the Gift-Eval test data Aksu et al. (2024) as well as in the Chronos pre-training data, which is used by several models including Tirex, Toto, and Sundial Ansari et al. (2024a); Auer et al. (2025); Liu et al. (2025); Cohen et al. (2025). As a precaution, Tirex and Toto excluded the dataset, though Tirex explicitly notes the difficulty of verifying whether the data is actually identical. Through our lineage analysis, we found that in the Chronos pre-training data, the "Solar" dataset includes solar power data from all US states at two frequencies: 5T (5 minutes) and 1H (hourly), where the 5T data was sourced directly from the original data provider [^5] https://www.nlr.gov/grid/solar-power-data. Gift-Eval, however, sourced only the Alabama subset as a test set, aggregated to 10T, 1H, 1D (daily), and 1W (weekly), through the LSTNet dataset repository [^6] https://github.com/laiguokun/multivariate-time-series-data/tree/master/solar-energy. LSTNet also sourced the data from the original provider but, as noted, included only Alabama rather than all states. In addition, the Alabama solar subset appears in the Monash Repository [^7] https://forecastingdata.org/ at frequencies 10T and 1W. This modified version of the source data was used as pretraining data by the models TinyTimeMixers and Lag-Llama Ekambaram et al. (2024); Rasul et al. (2024). While TinyTimeMixers and Lag-Llama were published before Gift-Eval, the authors of Sundial validate against Gift-Eval and state they excluded any overlaps without explicitly listing which datasets they excluded. This example illustrates how dataset provenance can become convoluted through multiple transformations and repositories, making it difficult for researchers to identify overlaps.
@@
While in the previous example the dataset name was constantly "Solar", relying on dataset names to avoid train-test overlaps is not sufficient. The identification of suitable datasets is complicated by dataset remixing and renaming. For example, the dataset "Elecdemand" from the Monash Repository is a scaled subset ($1/1000$) of the “Australian Electricity Demand” dataset. Similarly, the dataset "ElectricityLoadDiagrams20112014" appears under various names, such as "Electricity" in the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and "ECL" in the Informer (Zhou et al., 2021) collection. Dataset collections sometimes also contain different observations of the the same variables. For instance, the car ride shares dataset is included with same frequency and from the same year 2015 in both the Monash and GluonTS collections. Although both repositories cite the same GitHub source [^8] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they contain different time series of distinct For-Hire Vehicles (Uber in GluonTS vs Lyft in Monash). Consequently it is necessary to always compare the underlying data itself, keeping in mind it could have been transformed in various ways (scaling, handling missing values, renaming, resampling, etc.).
3.2 Indirect Information Leakage: Temporal Overlap of Correlated Train and Test Series
Direct overlap in data points between training and test sets is not the only cause for information leakage. In the context of time series data, a more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor and Narayanan, 2023). Often, these indirect dependencies stem from confounding bias, that is, an exogenous factor influencing observations from both datasets. For example, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains. However, not only global phenomena but also local factors can introduce such dependencies. Weather conditions at a specific location, for instance, directly influence solar energy production data, creating a causal relationship between meteorological measurements and power generation time series. Similarly, local traffic patterns may affect both air quality measurements and delivery time predictions within the same geographic area.
A key strength of TSFMs lies in their ability to learn subtle patterns in a flexible, data-driven manner and to apply these patterns to forecast future values of previously unseen time series. At large scale, however, TSFMs may inadvertently memorize patterns observed in one or more training time series and transfer them to series in the test set from the same time period. As long as training and test series are correctly ordered in time and do not overlap, this behavior does not constitute a serious issue. However, if training and test series overlap temporally and have a casual relationship (confounding or direct causation), such pattern transfer can result in information leakage. In the example illustrated in Figure 3, a highly flexible TSFM may learn the characteristic shape of the stock market decline caused by the global COVID-19 pandemic from the time series of the DAX index and reproduce it when tasked with forecasting the S&P 500 index over the same period.

### v3

# Rethinking Evaluation in the Era of Time Series Foundation Models: (Un)known Information Leakage Challenges

@@ Abstract
Time Series Foundation Models (TSFMs) represent a new paradigm for time-series forecasting, promising zero-shot predictions without the need for task-specific training or fine-tuning. However, similar to Large Language Models (LLMs), the evaluation of TSFMs is challenging: as training corpora grow increasingly large, it becomes difficult to ensure the integrity of the test sets used for benchmarking. An investigation of existing TSFM evaluation studies identifies two kinds of information leakage: (1) train-test sample overlaps arising from the multi-purpose reuse of datasets and (2) temporal overlap of correlated train and test series. Ignoring these forms of information leakage when benchmarking TSFMs risks producing overly optimistic performance estimates that fail to generalize to real-world settings. We therefore argue for the development of novel evaluation methodologies that avoid pitfalls already observed in both LLM and classical time-series benchmarking, and we call on the research community to adopt principled approaches to safeguard the integrity of TSFM evaluation.
@@ 1 Introduction
Time Series Foundation Models (TSFMs) represent an emerging paradigm in forecasting, drawing inspiration from the architecture and training methodologies of foundation models in natural language processing (NLP). In contrast to traditional time series models, TSFMs are pre-trained on large time series corpora, enabling zero-shot forecasting without task-specific adaptation (Liang et al., 2024). Over the last years, a highly dynamic landscape with a rapidly growing family of TSFMs has eveolved (Ansari et al., 2024; Auer et al., 2025; Cohen et al., 2025; Das et al., 2024).
@@
Yet, the very property that makes these TSFMs powerful, training globally on the world’s time series, creates structural evaluation problems, akin to issues recently observed in the evaluation of large language models (LLMs). In the NLP domain, training on vast portions of the internet has given rise to an “evaluation crisis" (Liao and Xiao, 2023), in which test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) and memorization effects (Chang et al., 2024) have led to overly optimistic performance estimates. As this perspective will show, current TSFM evaluations are vulnerable to analogous issues, leading to important implications for fair and reliable forecasting evaluation in the era of TSFMs.
While novel benchmark strategies such as benchmarking on held-out data (Aksu et al., 2024) and clean train-test splits (Qiu et al., 2024) have been proposed recently, these approaches cannot address the fundamental sources of information leakage in TSFMs. We identify two such sources (see Section 3): First, direct information leakage, arising from the multi-purpose reuse of public datasets across model training and evaluation pipelines. Second, indirect information leakage, arising from temporal overlap between correlated training and test series that often share a common causal driver — such as the COVID-19 pandemic simultaneously distorting correlated financial time series across geographies. Together, these sources risk turning current TSFM benchmarks into measures of memorization rather than generalization.
Our investigation reveals that information leakage is already present in TSFM benchmarking. Tracing the dataset lineage of 22 published TSFMs reveals no community consensus on which data should be used for training versus evaluation: in several documented cases, one model’s pre-training corpus is another model’s test set, making direct cross-model comparison nearly impossible. Yet, even benchmarks that carefully avoid such direct overlap remain exposed to indirect information leakage: a model trained on global stock indices through 2020 can learn the COVID-19 crash and exploit that structure when asked to forecast any correlated series from the same period, even one it has never directly seen (see Section 3).
The TSFM field risks repeating, in compressed time, the evaluation crisis that has undermined trust in LLM benchmarking. Preventing this doesn’t require incremental fixes to existing benchmarks, but a principled rethinking of what valid evaluation means when a model has been trained on the world’s time series.
2 Time Series Foundation Models: The Current State of Evaluation
Understanding the current state of TSFM evaluation requires first appreciating what distinguishes these models architecturally from classical forecasting methods, and how the benchmarking practices developed for those methods have been adapted to assess them.
@@ 2 Time Series Foundation Models: The Current State of Evaluation
2.1 Time Series Foundation Models: Global Training, Zero-Shot Inference
The foundation model paradigm, originally defined as training on broad data at scale to enable adaptation across downstream tasks (Bommasani et al., 2021), found its most prominent expression in LLMs such as the GPT and Llama families (OpenAI, 2023; Grattafiori et al., 2024), whose performance scales predictably with training data, parameters, and compute (Kaplan et al., 2020). Recognizing the structural sequence similarity between language modeling and time series analysis, recent work has adapted this paradigm to forecasting, producing a growing family of Transformer-based architectures that range from encoder-decoder designs (Ansari et al., 2024) to decoder-only (Das et al., 2024) and encoder-only models (Goswami et al., 2024). Since the emergence of TSFMs a few years ago, 22 models have been presented at top venues including TMLR, ICML, NeurIPS, ICLR, and WWW, as shown in Table 1.
@@ 2.1 Time Series Foundation Models: Global Training, Zero-Shot Inference
Chronos Ansari et al. (2024) 2024 TMLR
@@
Beyond the Transformer backbone, TSFMs draw on a range of architectural approaches: LLM-based models encode time series values directly as language tokens (Gruver et al., 2024; Xue and Salim, 2024), vision-based models represent them as grayscaled images (Chen et al., 2024), and reprogramming-based architectures mix textual and continuous representations (Jin et al., 2024). Despite this diversity, a key commonality runs through almost all TSFMs: pre-training on large, heterogeneous collections of time series that appear to follow scaling laws (Edwards et al., 2024), enabling zero-shot forecasting across a wide range of domains, frequencies, and horizons without retraining (MOMENT is the notable exception, requiring task-adaptive fine-tuning of its output layer (Goswami et al., 2024)). This zero-shot capability differs fundamentally from its NLP counterpart: while LLMs generalize to entirely different tasks (Brown et al., 2020), TSFMs generalize across domains, series, input lengths, and horizons, extending to new time series rather than new problem types (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024).
2.2 Classical and TSFM Benchmarking Practices
To rigorously assess a model’s predictive capabilities under evolving temporal dynamics, Time-Series Cross-Validation (TSCV) is widely regarded as best practice. Instead of using a single time-based split, TSCV uses the idea of rolling windows (see Figure 1 (a)). This evaluation strategy does not necessarily require retraining; instead, the training set can be fixed while the inference window is rolled forward, reducing evaluation costs to inference-only computation (Hewamalage et al., 2023).
Forecasting competitions, like the M-competitions or Kaggle, try to avoid the risk of test set contamination by soliciting forecasts for unpublished private test sets (Bojer and Meldgaard, 2021; Makridakis et al., 2022). These competitions are typically held at irregular intervals. The recent M6 competition took this further, introducing live evaluation using future values of financial assets as the target, bypassing reliance on historical data entirely (Makridakis et al., 2024).
While current TSFM benchmarking studies also rely on historical data for evaluation, their logic diverges from the above described time-series cross-validation strategy. Instead of using TSCV on one series as illustrated in Figure 1 (a), they typically evaluate models using many series but with only a single time-based split per series (Das et al., 2024; Gruver et al., 2024). Figure 1 (b) shows the focus of most TSFM evaluations. Here, the focus is not on temporal depth, but on generalization across heterogeneous series, especially in zero-shot scenarios where complete datasets or domains are held out of the pre-training (Shi et al., 2024; Ansari et al., 2024). This approach is focused on the generalization capabilities of TSFMs and prevents cherry-picking of test sets. Two prominent examples of this strategy are TSFM-Bench, which comprises 21 test sets with time series from diverse domains (e.g., energy, finance), statistical characteristics (e.g., stationary, trend), and frequencies (e.g., quarter-hourly, hourly) Li et al. (2025) and GIFT-Eval, which comprises 23 curated test sets with time series of seven domains, ten frequencies, and multivariate inputs (Aksu et al., 2024).
While this series cross validation focuses on evaluating the generalization capabilities of TSFMs, it is not without drawbacks. Roque et al. (2025) showed that the final ranking of models depends heavily on which series are selected as test sets. With just four cherry-picked test sets, out of the distribution of all available test sets, 46% of the benchmarked models could appear as state-of-the-art purely due to selection bias. Notably, deep learning models were found to be more susceptible to this variance than classical baselines. Especially when evaluating general-purpose TSFMs, generalization is a core evaluation focus.
Figure 1 (c) illustrates the current best practice in TSFM evaluation, that is, a combined time and series cross validation (Yu et al., 2016; Salinas et al., 2020). While this approach has first adopters, notably Goktas et al. (2025); Shchur et al. (2025), it is not yet commonly used in TSFM benchmarking.
2.3 Information Leakage in TSFM Benchmarking: A Recognized Challenge
As pre-training corpora grow larger and more heterogeneous, ensuring that evaluation data has not been seen during training becomes increasingly difficult. The same dynamic has been observed in LLM evaluation: models pre-trained on large internet crawls such as Common Crawl [^1] https://commoncrawl.org/ or The Pile [^2] https://pile.eleuther.ai/ had already seen many benchmark tasks during pre-training (Carlini et al., 2022; Chang et al., 2024), ultimately contributing to what has been described as an LLM “evaluation crisis” (Liao and Xiao, 2023). TSFMs face the same structural problems, and the community has begun to respond. As shown in the GIFT-Eval ablation studies, information leakage can strongly inflate evaluation performance (Aksu et al., 2024), and several recent benchmarks have developed explicit strategies to limit its impact (Aksu et al., 2024; Goktas et al., 2025; Xu et al., 2025). GIFT-Eval pairs an explicit pre-training corpus of approximately 230 billion data points with a dedicated test set of 177 million data points (Aksu et al., 2024). TempusBench draws on 48 forecasting tasks from sources currently absent from any known pre-training corpus, and further proposes live evaluation on real-world data (Goktas et al., 2025). Fidel-TS takes a similar approach through restricted live APIs (Xu et al., 2025).
Yet, none of these efforts have been preceded by a systematic analysis of where information leakage actually originates from, nor of how many distinct forms it takes. Without that foundation, benchmark design remains reactive rather than principled.
3 Two Distinct Pathways of Information Leakage in TSFM Evaluation
The integrity of any benchmark rests on a strict separation between training and evaluation data. For TSFMs, this separation is threatened along two distinct dimensions. The first is direct: datasets used for pre-training or fine-tuning reappear as test sets, undermining evaluation. The second is indirect and specific to time series: temporally overlapping but nominally disjoint series can be statistically correlated, allowing information about the test period to leak into the training signal without any test data point ever being seen directly. Evidence for both forms of leakage can already be found in current TSFM evaluations.
3.1 When Pre-Training Flexibility Comes at the Cost of Benchmark Comparability
Each TSFM team independently selects which datasets to use for pre-training, fine-tuning, and evaluation. Rightly so, since the freedom to train on diverse, large-scale corpora is central to the foundation model paradigm. Yet this very flexibility carries a structural side effect that has received little attention: once a dataset has been used for pre-training or fine-tuning by any model, it is no longer admissible as a foundation for benchmarking models against others. With each new model trained on a new combination of datasets, the set of valid datasets for benchmarking is reduced. A lineage analysis of 22 TSFMs published at leading machine learning venues (TMLR, ICML, NeurIPS, ICLR, WWW) and their workshops through January 2026 (Table 1) makes this concrete. For each model, every dataset was categorized into one of three roles: pre-training (used only for initial model training and thus excluded from evaluation), train/test (split temporally for fine-tuning and in-domain evaluation), or zero-shot (reserved exclusively for held-out evaluation on unseen data).
@@ 3.1 When Pre-Training Flexibility Comes at the Cost of Benchmark Comparability
The lineage map (Figure 2; full dataset-level analysis in Section 6) confirms this: across 401 datasets identified, each TSFM team assembled a distinct combination of pre-training, train/test, and zero-shot datasets, and no single dataset has been universally reserved for evaluation. Only 6% of these 401 datasets have never appeared in any model’s pre-training or fine-tuning corpus. These are the only candidates that could, in principle, support genuine zero-shot comparisons across the full set of published TSFMs.
The pervasiveness of this problem is best illustrated through concrete examples. The Australian Electricity Demand dataset from the Monash collection (Godahewa et al., 2021) has been used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM), making it practically useless for any cross-model comparison. Similarly, the Informer collection (Zhou et al., 2021), widely treated as a zero-shot benchmark, contains individual series such as ETTh1, ETTh2, and ETTm1 that have also been used for pre-training (e.g., in Lag-Llama) or train/test evaluation (e.g., in UniTime).
@@
These overlaps are often difficult to detect, because datasets are routinely remixed, renamed, and redistributed across repositories. For instance, the dataset “Elecdemand” from the Monash Repository is a scaled subset ($1/1000$) of “Australian Electricity Demand”. The dataset “ElectricityLoadDiagrams20112014” appears under the name “Electricity” in both the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and as “ECL” in the Informer collection (Zhou et al., 2021). Even datasets sharing the same name may contain different underlying data: the car ride shares dataset is included from the same year 2015 in both the Monash and GluonTS collections. Although both repositories cite the same GitHub source [^3] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they contain different time series of distinct For-Hire Vehicles (Uber in GluonTS vs. Lyft in Monash). Consequently, it is necessary to always compare the underlying data itself, keeping in mind that it could have been transformed in various ways (scaling, handling missing values, renaming, resampling, etc.).
A particularly telling example of how dataset provenance can become convoluted through multiple transformations and repositories involves the “Solar” dataset. The Chronos pre-training corpus includes solar power data from all US states at two frequencies: 5T (5 minutes) and 1H (hourly), sourced directly from the original data provider [^4] https://www.nlr.gov/grid/solar-power-data. GIFT-Eval, however, includes only the Alabama subset as a test set, aggregated to 10T, 1H, 1D (daily), and 1W (weekly), sourced through the LSTNet dataset repository [^5] https://github.com/laiguokun/multivariate-time-series-data/tree/master/solar-energy. LSTNet also sourced the data from the original provider but, as noted, included only Alabama rather than all states. In addition, the Alabama solar subset appears in the Monash Repository [^6] https://forecastingdata.org/ at frequencies 10T and 1W. This modified version was used as pretraining data by TinyTimeMixers and Lag-Llama (Ekambaram et al., 2024; Rasul et al., 2024). While TinyTimeMixers and Lag-Llama were published before GIFT-Eval, the authors of Sundial validate against GIFT-Eval and state they excluded any overlaps without explicitly listing which datasets they excluded (Aksu et al., 2024; Liu et al., 2025). The authors of TiRex and Toto discovered this entanglement and excluded the dataset as a precaution, though TiRex explicitly notes the difficulty of verifying whether the data across repositories is actually identical (Auer et al., 2025; Cohen et al., 2025).
The consequences of undetected contamination can be severe. As summarized in Supplementary Information Section A, documented cases of leakage between training and test sets have led to over 50% better test scores. Several such cases suggest that overlaps between train and test data are not always identified during the peer-review process (Saravanan et al., 2024; Montet et al., 2025; Li et al., 2024b). This is not surprising: exactly determining on which datasets a model has been trained requires careful reading of the original paper, its appendix, and sometimes even analyzing published source code. During the lineage analysis, a potential overlap surfaced between the pre-training and test set in TimesFM (Das et al., 2024), as the traffic hourly dataset appears to have been used during pretraining while also being included in the reported Monash evaluation benchmark. This anecdotal example underscores how difficult it is to maintain a reliable overview of datasets; a challenge that only intensifies as the number of published TSFMs grows.
3.2 When Temporal Overlap Turns Correlation Into Leakage
Time series carry a special structural property: temporal correlation between nominally independent series. This correlation often has causal roots, yet TSFMs do not learn causal relationships; they learn from observed statistical patterns alone. When an exogenous shock affects multiple domains simultaneously — a pandemic, a financial crisis, a geopolitical shift — it imprints a recognizable signature across many series at once. The same effect operates at a local scale: weather conditions at a specific location directly shape solar energy production (Aryandoust et al., 2022), and local traffic patterns simultaneously influence air quality measurements and delivery time predictions. A model trained on any of these series during the affected period has, in effect, learned something about all of them, even those it never encountered directly. While spurious correlation without any shared cause could in principle produce a similar effect, it may also lead to worse predictions, as the model transfers patterns that bear no true relationship to the test series. But the systematic correlation arising from genuine common drivers is the more prevalent and serious concern.